# Multivariate Probabilistic Time Series Forecasting with Correlated Errors

### v3

@@ 4 Our Method
Our methodology builds upon the formulation outlined in Eq. (2), employing an autoregressive model as its foundational framework. Using RNN as an example, a probabilistic forecasting model consists of two components. Firstly, it incorporates a transition model $f_{\Theta}$ to capture the dynamics of state transitions $\mathbf{h}_{t}=f_{\Theta}\left(\mathbf{h}_{t-1},\mathbf{z}_{t-1},\mathbf{x}_{t}\right)$, thus inherently having autoregressive properties. Secondly, it integrates a distribution head, represented by $\theta$, which is tasked with mapping $\mathbf{h}_{t}$ to the parameters governing the desired probability distribution. Following GPVar [2], our approach employs the multivariate Gaussian distribution as the distribution head. The time series variable can be decomposed into a deterministic mean component and a random error component $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$, where $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. To efficiently model covariance $\boldsymbol{\Sigma}_{t}$ for large $N$, GPVar adopts a low-rank-plus-diagonal parameterization $\boldsymbol{\Sigma}_{t}=\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}+\operatorname{diag}{(\mathbf{d}_{t})}$, where $\boldsymbol{L}_{t}\in\mathbb{R}^{B\times R}$ ($R\ll B)$ and $\boldsymbol{d}_{t}\in\mathbb{R}_{+}^{B}$. Autoregressive models based on Gaussian likelihood are usually built on the assumption that $\boldsymbol{\eta}_{t}$ are independently distributed following a multivariate Gaussian distribution. The log-likelihood of the distribution serves as the loss function for optimizing the model:
@@
The parameters $\theta(\mathbf{h}_{t})$ are parameterized as $(\boldsymbol{\mu}_{t},\boldsymbol{L}_{t},\mathbf{d}_{t})$, where $\boldsymbol{\mu}_{t}\in\mathbb{R}^{N}$ represents the mean vector of the distribution. $\boldsymbol{L}_{t}\in\mathbb{R}^{N\times R}$ and $\mathbf{d}_{t}\in\mathbb{R}_{+}^{N}$ correspond to the covariance factor and diagonal elements in the low-rank parameterization of the multivariate Gaussian distribution. We use shared mapping functions for all time series:
@@
where $\mathbf{h}_{i,t}\in\mathbb{R}^{H\times 1}$, $\mathbf{w}_{\mu}\in\mathbb{R}^{H\times 1}$, $\mathbf{w}_{d}\in\mathbb{R}^{H\times 1}$, and $W_{l}\in\mathbb{R}^{R\times H}$ are parameters. Since the parameters of mapping functions are shared for all time series, we can view $\mathbf{z}_{t}$ as a Gaussian process assessed at points $\mathbf{h}_{i,t}$. Consequently, we can leverage a random subset of time series to compute the Gaussian likelihood-based loss in each iteration. In other words, we can train the model with a substantially reduced batch size $B<N$.
@@ 4.1 Training with Correlated Errors
We expand upon the approach introduced in [8] to handle cross-correlated errors in a multivariate context by introducing time-dependent error terms $\boldsymbol{\eta}_{t}$ into the GLS loss. In many existing deep probabilistic forecasting models, such as GPVar [2], a training batch typically consists of a sample slice of $B$ time series spanning a temporal length of $P+Q$, where $P$ is the conditioning range and $Q$ is the prediction range. Gaussian likelihood is evaluated independently at each time step within the prediction range through one-step-ahead prediction. However, this direct approach overlooks the serial correlation of errors across consecutive time steps. To address this drawback, we propose a modification to the likelihood function by introducing a dynamic covariance that accommodates the temporal dependence of the error term, as illustrated in Fig. 2. To achieve this, we organize $D$ smaller slices of time series with a temporal length of $P+1$ (i.e., $Q=1$), sorted by the prediction start time in sequential order, where $D$ represents the time horizon in which we consider cross-correlation. It is evident that the new batch structure is a reconstruction of the conventional training batch, covering the same time horizon if $D=Q$. An example of the collection of target time series variables in a batch covering cross-correlation horizon $D$ is given by
@@
where for time point $t^{\prime}$, $\boldsymbol{\mu}_{t^{\prime}}$, $\boldsymbol{L}_{t^{\prime}}$ and $\mathbf{d}_{t^{\prime}}$ are the output of the model. The covariance parameterization in GPVar corresponds to
@@
where $\mathbf{r}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{R})$ is a low-dimensional latent variable, and $\boldsymbol{\varepsilon}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\operatorname{diag}{(\mathbf{d}_{t^{\prime}})})$ is an additional error that is independent of $\mathbf{r}_{t^{\prime}}$. We denote by $\boldsymbol{z}_{t}^{\text{bat}}=\operatorname{vec}\left(\mathbf{z}_{t-D+1:t}\right)\in\mathbb{R}^{DB}$ the collection of target time series variables in a batch, where $\operatorname{vec}(\cdot)$ is an operator that stacks all the columns of a matrix into a vector. Similarly, we have $\boldsymbol{\mu}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{r}_{t}^{\text{bat}}\in\mathbb{R}^{DR}$, $\boldsymbol{\varepsilon}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{d}_{t}^{\text{bat}}\in\mathbb{R}_{+}^{DB}$, and $\boldsymbol{L}_{t}^{\text{bat}}=\operatorname{blkdiag}(\{\boldsymbol{L}_{t^{\prime}}\}_{t^{\prime}=t-D+1}^{t})\in\mathbb{R}^{DB\times DR}$ has a block diagonal structure (see Fig. 2). We then have the batch-wise decomposition:
@@
The default GPVar model assumes the latent variable $\boldsymbol{r}_{t}$ to be temporally independent, i.e., $\operatorname{Cov}\left(\boldsymbol{r}_{s},\boldsymbol{r}_{t}\right)=\boldsymbol{0},\forall s\neq t$. However, this assumption cannot accommodate the potential cross-correlation in the error. We next introduce temporal dependencies in the latent variable in a batch by assuming $\boldsymbol{r}_{t}^{\text{bat}}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}\right)$, where $\boldsymbol{C}_{t}$ is a $D\times D$ dynamic correlation matrix. This approach corresponds to assuming that the rows in the matrix $\mathbf{r}_{t-D+1:t}=\left[\mathbf{r}_{t-D+1},\ldots,\mathbf{r}_{t}\right]$ are independent and identically distributed following $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\right)$. To efficiently capture dynamic patterns over time, we follow Zheng et al. [7] and express $\boldsymbol{C}_{t}$ as a dynamic weighted sum base kernel matrices: $\boldsymbol{C}_{t}=\sum_{m=1}^{M}w_{m,t}\boldsymbol{K}_{m}$, where $w_{m,t}\geq 0$ (with $\sum_{m}w_{m,t}=1$) represents the weights for each component. For simplicity, we model each component $\boldsymbol{K}_{m}$ using a kernel matrix generated from a squared-exponential (SE) kernel function, with the $(i,j)$th entry being $\boldsymbol{K}^{ij}_{m}=\exp(-\frac{(i-j)^{2}}{l_{m}^{2}})$, with different lengthscales $l$ (e.g., $l=1,2,3,\dots$). In addition, we incorporate an identity matrix into the additive structure to account for the independent noise process. This parameterization ensures that $\boldsymbol{C}_{t}$ is a positive definite symmetric matrix with unit diagonals, making it a valid correlation matrix. The weights for these components are derived from the hidden state $\mathbf{h}_{t}$ at each time step through a small neural network, with the number of nodes in the output layer set to $M$ (i.e., the number of components). A softmax layer is then used to ensure that these weights are summed up to 1. Note that the parameters of this network will be learned simultaneously with those of the base model.
@@
It is not difficult to derive that for any $i,j\in\{0,1,\ldots,D-1\}$ and $i\neq j$, the proposed model can create cross-covariance $\operatorname{Cov}\left(\boldsymbol{\eta}_{t-i},\boldsymbol{\eta}_{t-j}\right)=\boldsymbol{C}_{t}^{ij}\boldsymbol{L}_{t-i}\boldsymbol{L}_{t-j}^{\top}$ for times $t-i$ and $t-j$, which is no longer $\boldsymbol{0}$. In this sense, our model can better characterize the empirical cross-covariance matrices of the residuals (see empirical examples in Fig. 1). As $\boldsymbol{\mu}_{t}^{\text{bat}}$, $\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{d}_{t}^{\text{bat}}$ are default outputs of the base probabilistic model, we can compute the overall likelihood (with overlapped data) by
@@
Here, computing the log-likelihood involves evaluating the inverse and the determinant of $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ with size $DB\times DB$, for which a naïve implementation has a prohibitive time complexity of $\mathcal{O}\left(D^{3}B^{3}\right)$. However, it should be noted that the covariance $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ in our parameterization is of the form $\boldsymbol{E}+\boldsymbol{A}\boldsymbol{C}\boldsymbol{A}^{\top}$, where $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d}_{t}^{\text{bat}})$ is diagonal, $\boldsymbol{A}=\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{C}=\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$. This allows us to leverage the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma to simplify the computation:
@@
Then, the likelihood calculation only requires computing the inverse and determinant of a $DR\times DR$ matrix, specifically $\boldsymbol{C}^{-1}+\boldsymbol{A}^{\top}\boldsymbol{E}^{-1}\boldsymbol{A}$. Both of these computations can be efficiently performed using its Cholesky factorization. Detailed computations are provided in Appendix §A.2.
Modeling the latent process $\boldsymbol{r}_{t}$ offers several advantages. Firstly, because $\boldsymbol{r}_{t}$ has a much lower dimension than $\boldsymbol{\varepsilon}_{t}$, modeling the cross-correlation of $\boldsymbol{r}_{t}$ results in a significantly smaller $DR\times DR$ covariance matrix compared to the $DB\times DB$ covariance matrix of $\boldsymbol{\varepsilon}_{t}$. Secondly, since $\boldsymbol{r}_{t}$ follows an isotropic Gaussian distribution, the covariance of $\boldsymbol{r}_{t}^{\text{bat}}$ can be parameterized with a Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$. This greatly simplifies the task into learning a $D\times D$ correlation matrix shared by all time series in a batch. Lastly, similar to GPVar, we can still train the model using a subset of time series in each iteration, maintaining computational efficiency.
@@ 4.2 Multistep-ahead Rolling Prediction
Autoregressive models conduct multistep-ahead forecasting in an iterative manner—the model generates a sample at each time step during prediction, serving as input for the subsequent step, and the process continues until the desired prediction range is reached. Our approach enhances this process in a similar way as in Zheng et al. [8] by offering additional calibration based on the learned correlation matrix $\boldsymbol{C}_{t}$. Assuming observations are available up to time step $t$, the conditional distribution of $\boldsymbol{\eta}_{t+1}$ given errors in the past $(D-1)$ steps, can be derived as follows:
@@
where $\boldsymbol{\eta}_{\text{obs}}=\operatorname{vec}\left(\left[\boldsymbol{\eta}_{t-D+2},\ldots,\boldsymbol{\eta}_{t-1},\boldsymbol{\eta}_{t}\right]\right)\in\mathbb{R}^{(D-1)B}$ represents the set of residuals, accessible at forecasting step $t+1$. Here, $\boldsymbol{\Sigma}_{\text{obs}}$ is a $(D-1)B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{\text{obs}}$, and $\boldsymbol{\Sigma}_{*}$ is a $B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{t+1}$ and $\boldsymbol{\eta}_{\text{obs}}$, i.e., $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}=\begin{bmatrix}\boldsymbol{\Sigma}_{\text{obs}}&\boldsymbol{\Sigma}_{*}^{\top}\\ \boldsymbol{\Sigma}_{*}&\boldsymbol{\Sigma}_{t+1}\end{bmatrix}$. For conciseness, we omit the time index $t$ in $\boldsymbol{\Sigma}_{\text{obs}}$, $\boldsymbol{\Sigma}_{*}$ and $\boldsymbol{\eta}_{\text{obs}}$. Given that $\boldsymbol{\mu}_{t+1}$ is a deterministic output from the base model, a sample of the target variables $\tilde{\mathbf{z}}_{t+1}$ can be derived by initially drawing a sample $\tilde{\boldsymbol{\eta}}_{t+1}$ from Eq. (12) and then combining it with the predicted mean vector $\boldsymbol{\mu}_{t+1}$ by $\tilde{\mathbf{z}}_{t+1}=\boldsymbol{\mu}_{t+1}+\tilde{\boldsymbol{\eta}}_{t+1}$.
Taking the sample $\tilde{\boldsymbol{\eta}}_{t+1}$ as an observed residual, we can iteratively execute the process described in Eq. (12) to derive a trajectory of $\{\tilde{\mathbf{z}}_{t+q}\}_{q=1}^{Q}$. By repeating this procedure, we can generate a certain number of samples to characterize the predictive distribution at each time step.

### v4

@@ 4 Our Method
Our methodology builds upon the formulation outlined in Eq. (2), employing an autoregressive model as its foundational framework. Using an RNN as an example, a probabilistic forecasting model consists of two components. Firstly, it incorporates a transition model $f_{\Theta}$ to capture the dynamics of state transitions $\mathbf{h}_{t}=f_{\Theta}\left(\mathbf{h}_{t-1},\mathbf{z}_{t-1},\mathbf{x}_{t}\right)$, thus inherently having autoregressive properties. Second, it integrates a distribution head, represented by $\theta$, which maps $\mathbf{h}_{t}$ to the parameters of the desired probability distribution. Following GPVar [3], our approach employs the multivariate Gaussian distribution as the distribution head. The time series variable can be decomposed into a deterministic mean component and a random error component $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$, where $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. To efficiently model the covariance $\boldsymbol{\Sigma}_{t}$ for large $N$, GPVar adopts a low-rank-plus-diagonal parameterization $\boldsymbol{\Sigma}_{t}=\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}+\operatorname{diag}{(\mathbf{d}_{t})}$, where $\boldsymbol{L}_{t}\in\mathbb{R}^{N\times R}$ ($R\ll N)$ and $\boldsymbol{d}_{t}\in\mathbb{R}_{+}^{N}$. Autoregressive models based on Gaussian likelihood typically assume that $\boldsymbol{\eta}_{t}$ are independently distributed following a multivariate Gaussian distribution. The log-likelihood of the distribution serves as the loss function for optimizing the model:
@@
The parameters $\theta(\mathbf{h}_{t})$ are parameterized as $(\boldsymbol{\mu}_{t},\boldsymbol{L}_{t},\mathbf{d}_{t})$, where $\boldsymbol{\mu}_{t}\in\mathbb{R}^{N}$ represents the mean vector of the distribution. $\boldsymbol{L}_{t}$ and $\mathbf{d}_{t}$ correspond to the covariance factor and diagonal elements in the low-rank parameterization of the multivariate Gaussian distribution. We use shared mapping functions for all time series:
@@
where $\mathbf{h}_{i,t}\in\mathbb{R}^{H}$, $\mathbf{w}_{\mu}\in\mathbb{R}^{H}$, $\mathbf{w}_{d}\in\mathbb{R}^{H}$, and $W_{l}\in\mathbb{R}^{R\times H}$ are parameters. Since the parameters of the mapping functions are shared across all time series, we can use a random subset of time series to compute the Gaussian likelihood-based loss in each optimization step, as any subset of $\mathbf{z}_{t}$ will still follow a multivariate Gaussian distribution. In other words, we can train the model with a substantially reduced batch size $B<N$.
@@ 4.1 Training with Correlated Errors
We build upon the approach introduced in [9] to address cross-correlated errors in a multivariate context by introducing time-dependent error terms $\boldsymbol{\eta}_{t}$ into the GLS loss. In many existing deep probabilistic forecasting models, such as GPVar [3], a training batch typically consists of a sample slice of $B$ time series spanning a temporal length of $P+Q$, where $P$ is the conditioning range and $Q$ is the prediction range. The Gaussian likelihood is evaluated independently at each time step within the prediction range through one-step-ahead predictions. However, this approach overlooks the serial correlation of errors across consecutive time steps. To address this limitation, we propose modifying the likelihood function by introducing a dynamic covariance that accommodates the temporal dependence of the error term, as illustrated in Fig. 2. To achieve this, we organize $D$ smaller slices of time series with a temporal length of $P+1$ (i.e., $Q=1$), sorted by the prediction start time in sequential order, where $D$ represents the time horizon over which we consider cross-correlation. The new batch structure effectively reconstructs the conventional training batch, covering the same time horizon when $D=Q$. An example of the collection of target time series variables in a batch covering cross-correlation horizon $D$ is given by
@@
where for time point $t^{\prime}$, $\boldsymbol{\mu}_{t^{\prime}}$, $\boldsymbol{L}_{t^{\prime}}$ and $\mathbf{d}_{t^{\prime}}$ are the outputs of the model. The covariance parameterization in GPVar corresponds to
@@
where $\mathbf{r}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{R})$ is a low-dimensional latent variable, and $\boldsymbol{\varepsilon}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\operatorname{diag}{(\mathbf{d}_{t^{\prime}})})$ is an additional error independent of $\mathbf{r}_{t^{\prime}}$. We denote $\boldsymbol{z}_{t}^{\text{bat}}=\operatorname{vec}\left(\mathbf{z}_{t-D+1:t}\right)\in\mathbb{R}^{DB}$ as the collection of target time series variables in a batch, where $\operatorname{vec}(\cdot)$ is an operator that stacks all the columns of a matrix into a vector. Similarly, we define $\boldsymbol{\mu}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{r}_{t}^{\text{bat}}\in\mathbb{R}^{DR}$, $\boldsymbol{\varepsilon}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{d}_{t}^{\text{bat}}\in\mathbb{R}_{+}^{DB}$, and $\boldsymbol{L}_{t}^{\text{bat}}=\operatorname{blkdiag}(\{\boldsymbol{L}_{t^{\prime}}\}_{t^{\prime}=t-D+1}^{t})\in\mathbb{R}^{DB\times DR}$, where $\boldsymbol{L}_{t}^{\text{bat}}$ has a block diagonal structure (see Fig. 2). The batch-wise decomposition is then expressed as
@@
The default GPVar model assumes the latent variable $\boldsymbol{r}_{t}$ is temporally independent, meaning $\operatorname{Cov}\left(\boldsymbol{r}_{s},\boldsymbol{r}_{t}\right)=\boldsymbol{0},\forall s\neq t$. However, this assumption cannot capture the potential cross-correlation in the errors. To address this, we introduce temporal dependencies in the latent variable within a batch by assuming $\boldsymbol{r}_{t}^{\text{bat}}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}\right)$, where $\boldsymbol{C}_{t}$ is a dynamic $D\times D$ correlation matrix. This approach assumes that the rows in the matrix $\mathbf{r}_{t-D+1:t}=\left[\mathbf{r}_{t-D+1},\ldots,\mathbf{r}_{t}\right]$ are independent and identically distributed, following $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\right)$. To efficiently capture dynamic patterns over time, we follow Zheng et al. [9] and express $\boldsymbol{C}_{t}$ as a dynamic weighted sum base kernel matrices: $\boldsymbol{C}_{t}=\sum_{m=1}^{M}w_{m,t}\boldsymbol{K}_{m}$, where $w_{m,t}\geq 0$ (with $\sum_{m}w_{m,t}=1$) represents the weights for each component. For simplicity, we model each component $\boldsymbol{K}_{m}$ using a kernel matrix generated from a squared-exponential (SE) kernel function, where the $(i,j)$-th entry is $\boldsymbol{K}^{ij}_{m}=\exp(-\frac{(i-j)^{2}}{l_{m}^{2}})$, with different lengthscales $l_{m}$ (e.g., $l=1,2,3,\dots$). In addition, we incorporate an identity matrix into the additive structure to account for the independent noise process. This parameterization ensures that $\boldsymbol{C}_{t}$ is a positive definite symmetric matrix with unit diagonals, making it a valid correlation matrix. The weights for these components are derived from the hidden state $\mathbf{h}_{t}$ at each time step through a small neural network, with the number of nodes in the output layer set to $M$ (i.e., the number of components). A softmax layer is used to ensure that these weights are summed up to 1. Note that the parameters of this network will be learned simultaneously with those of the base model.
Figure 3: Illustration of the training process. Following [3], time series dimensions are randomly sampled, and the base model (e.g., RNNs) is unrolled for each dimension individually (e.g., 1, 2, 4, followed by 1, 3, 4 as depicted). The model parameters are shared across all time series dimensions. A batch of time series variables $\boldsymbol{z}_{t}^{\text{bat}}$ contains time series vectors $\boldsymbol{z}_{t}$ covering time steps from $t-D+1$ to $t$. In contrast to [3], our approach explicitly models dependencies over the extended temporal window from $t-D+1$ to $t$ during training.
@@
It is straightforward to derive that for any $i,j\in\{0,1,\ldots,D-1\}$ and $i\neq j$, the proposed model creates cross-covariance $\operatorname{Cov}\left(\boldsymbol{\eta}_{t-i},\boldsymbol{\eta}_{t-j}\right)=\boldsymbol{C}_{t}^{ij}\boldsymbol{L}_{t-i}\boldsymbol{L}_{t-j}^{\top}$ between times $t-i$ and $t-j$, which is no longer $\boldsymbol{0}$. While this parameterization results in a non-stationary multivariate process through varying coregionalization [22, 23], a key difference is that both the coregionalization coefficient matrix $\boldsymbol{L}_{t}$ and the temporal correlation $\boldsymbol{C}_{t}$ are generated by a deep neural network. In this sense, our model can better characterize the empirical cross-covariance matrices of the residuals (see empirical examples in Fig. 1). As $\boldsymbol{\mu}_{t}^{\text{bat}}$, $\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{d}_{t}^{\text{bat}}$ are default outputs of the base probabilistic model, we can compute the overall likelihood (with overlapped data) as
@@
Here, computing the log-likelihood involves evaluating the inverse and the determinant of $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ with size $DB\times DB$, for which a naive implementation has a prohibitive time complexity of $\mathcal{O}\left(D^{3}B^{3}\right)$. However, our parameterization of $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ as $\boldsymbol{E}+\boldsymbol{A}\boldsymbol{C}\boldsymbol{A}^{\top}$, where $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d}_{t}^{\text{bat}})$, $\boldsymbol{A}=\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{C}=\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$, allows us to leverage the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma to simplify the computation:
@@
Then, the likelihood calculation only requires computing the inverse and determinant of a $DR\times DR$ matrix, specifically $\boldsymbol{C}^{-1}+\boldsymbol{A}^{\top}\boldsymbol{E}^{-1}\boldsymbol{A}$. These computations can be efficiently performed using Cholesky factorization. Detailed computations are provided in Appendix §A.2.
Modeling the latent process $\boldsymbol{r}_{t}$ offers several advantages. Firstly, because $\boldsymbol{r}_{t}$ has a much lower dimension than $\boldsymbol{\varepsilon}_{t}$, modeling the cross-correlation of $\boldsymbol{r}_{t}$ results in a significantly smaller $DR\times DR$ covariance matrix compared to the $DB\times DB$ covariance matrix of $\boldsymbol{\varepsilon}_{t}$. Secondly, since $\boldsymbol{r}_{t}$ follows an isotropic Gaussian distribution, the covariance of $\boldsymbol{r}_{t}^{\text{bat}}$ can be parameterized with a Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$. This greatly simplifies the task into learning a $D\times D$ correlation matrix shared by all time series in a batch. Lastly, similar to GPVar, we can still train the model in an end-to-end manner using a subset of time series in each iteration to ensure computational efficiency (Fig. 3).
@@ 4.2 Multistep-ahead Rolling Prediction
Autoregressive models perform multistep-ahead forecasting in an iterative manner, where the model generates a sample at each time step during prediction, using it as input for the subsequent step, and continuing this process until the desired prediction range is reached. Our approach enhances this process, similar to Zheng et al. [9], by offering additional calibration based on the learned correlation matrix $\boldsymbol{C}_{t}$. Assuming observations are available up to time step $t$, the conditional distribution of $\boldsymbol{\eta}_{t+1}$ given errors in the past $(D-1)$ steps, can be derived as
@@
where $\boldsymbol{\eta}_{\text{obs}}=\operatorname{vec}\left(\left[\boldsymbol{\eta}_{t-D+2},\ldots,\boldsymbol{\eta}_{t-1},\boldsymbol{\eta}_{t}\right]\right)\in\mathbb{R}^{(D-1)B}$ represents the set of residuals, accessible at forecasting step $t+1$. Here, $\boldsymbol{\Sigma}_{\text{obs}}$ is a $(D-1)B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that captures the covariance of $\boldsymbol{\eta}_{\text{obs}}$, and $\boldsymbol{\Sigma}_{*}$ is a $B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ representing the covariance between $\boldsymbol{\eta}_{t+1}$ and $\boldsymbol{\eta}_{\text{obs}}$, i.e., $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}=\begin{bmatrix}\boldsymbol{\Sigma}_{\text{obs}}&\boldsymbol{\Sigma}_{*}^{\top}\\ \boldsymbol{\Sigma}_{*}&\boldsymbol{\Sigma}_{t+1}\end{bmatrix}$. For conciseness, we omit the time index $t$ in $\boldsymbol{\Sigma}_{\text{obs}}$, $\boldsymbol{\Sigma}_{*}$ and $\boldsymbol{\eta}_{\text{obs}}$. Since $\boldsymbol{\mu}_{t+1}$ is a deterministic output from the base model, a sample of the target variables $\tilde{\mathbf{z}}_{t+1}$ can be derived by first drawing a sample $\tilde{\boldsymbol{\eta}}_{t+1}$ from Eq. (12), then combining it with the predicted mean vector $\boldsymbol{\mu}_{t+1}$ as $\tilde{\mathbf{z}}_{t+1}=\boldsymbol{\mu}_{t+1}+\tilde{\boldsymbol{\eta}}_{t+1}$. It should be noted that we can still leverage the Sherman-Morrison-Woodbury identity when computing the inverse $\boldsymbol{\Sigma}_{\text{obs}}^{-1}$.
By taking the sample $\tilde{\boldsymbol{\eta}}_{t+1}$ as an observed residual, we can iteratively apply the process described in Eq. (12) to derive a trajectory of $\{\tilde{\mathbf{z}}_{t+q}\}_{q=1}^{Q}$. Repeating this procedure allows us to generate multiple samples, characterizing the predictive distribution at each time step.